# Organizing Unstructured Image Collections using Natural Language

### v1

@@ Abstract
Organizing unstructured visual data into semantic clusters is a key challenge in computer vision. Traditional deep clustering (DC) approaches focus on a single partition of data, while multiple clustering (MC) methods address this limitation by uncovering distinct clustering solutions. The rise of large language models (LLMs) and multimodal LLMs (MLLMs) has enhanced MC by allowing users to define clustering criteria in natural language. However, manually specifying criteria for large datasets is impractical. In this work, we introduce the task Semantic Multiple Clustering (SMC) that aims to automatically discover clustering criteria from large image collections, uncovering interpretable substructures without requiring human input. Our framework, Text Driven Semantic Multiple Clustering (TeDeSC), uses text as a proxy to concurrently reason over large image collections, discover partitioning criteria, expressed in natural language, and reveal semantic substructures. To evaluate TeDeSC, we introduce the COCO-4c and Food-4c benchmarks, each containing four grouping criteria and ground-truth annotations. We apply TeDeSC to various applications, such as discovering biases and analyzing social media image popularity, demonstrating its utility as a tool for automatically organizing image collections and revealing novel insights.
Disclaimer: Potentially sensitive content.
Figure 1: Top: Semantic Multiple Clustering (SMC) deals with automatically organizing an unstructured image collection into semantically meaningful and human interpretable groups (or semantic clusters), under multiple shared themes (or clustering criteria), without requiring any prior information. Bottom: Our proposed SMC system enables various applications like discovering novel biases in text-to-image generative (T2I) models.
@@ 1 Introduction
Organizing large volumes of unstructured visual data into semantic clusters is a crucial problem in computer vision and has traditionally been addressed through the lens of unsupervised deep clustering (DC) (Caron et al., 2018; Van Gansbeke et al., 2020). Despite the advancements and improved performance, DC remains limited due to its inherently ill-posed nature (Estivill-Castro, 2002), as a dataset can be clustered in multiple ways (e.g., a dataset of a standard deck of cards can be grouped by suit or rank). Furthermore, the clusters discovered by DC rely on factors such as the inductive biases of the network, augmentations, and pretext tasks; and do not necessarily adhere to the a particular partitioning criteria that a user may have in mind (e.g., suit).
The limitation of DC in exploring visual data only along a single partition is addressed by multiple clustering (MC) methods (Yu et al., 2024), which uncover distinct clustering solutions. Multiple partitions reveal hidden patterns in the data and allow for analysis from different perspectives. The recent advent of large language models (LLMs) and multimodal LLMs (MLLMs) has further enhanced MC by enabling users to specify the clustering criterion (e.g., group by suit) in natural language (Kwon et al., 2024; Yao et al., 2024). While incorporating a human in the loop is appealing, as it grants the user control over the clustering results, specifying a meaningful text criterion still requires prior understanding of the image collections. This task becomes cumbersome, especially due to the proliferation and high scene complexity inherent in large volumes of visual data.
@@
Prior User-provided Criteria
Knowledge # Clusters
Out. Multiple Clustering
Interpretability
In this work, we propose to ease the burden of specifying text criteria off the user by automatically discovering the clustering criteria, and introduce the task of Semantic Multiple Clustering (SMC), as shown in Fig. 1(top). To tackle SMC, we propose a general two-stage framework Text Driven Semantic Multiple Clustering (TeDeSC), powered by cutting-edge MLLMs and LLMs, that first discovers latent criteria (e.g., Activity and Location) from a given collection of images, and then groups the images into semantic clusters (e.g., “Surfing”,“Skateboarding” under Activity) for every discovered clustering criterion. To reason at a dataset level, TeDeSC first translates the visual data into textual data (e.g., captions), which in turn are then collectively used as a proxy for the LLM to discover the hidden patterns in the image collection. The uniqueness of TeDeSC lies in its: (i) comprehensiveness, as it can exhaustively discover multiple clustering criteria that may not be evident to a user; (ii) interpretability, as it outputs cluster names in natural language as opposed to traditional MC methods; (iii) flexibility, as it can discover clusters at multiple granularities (coarse to fine-grained), without needing to specify the number of clusters a priori; and (iv) generality, as, unlike the existing work (Yao et al., 2024), it is not limited to object-centric images but can handle complex scenes having fine-grained details. We summarize the key differences among various clustering solutions in Tab. 1 and elaborate in App. A.
Current benchmarks for evaluating multiple clustering either lack realism (Clevr-4c, Fruit-2c, Card-2c) or offer a limited number of criteria (Action-3c), as shown in Fig. 2. To advance research in SMC, we introduce two challenging new benchmarks, COCO-4c and Food-4c, which depict images in daily life contexts and allow the dataset to be clustered as per four distinct criteria. Extensive experimental analyses on both existing and novel benchmarks demonstrate that the proposed TeDeSC can effectively discover meaningful clustering criteria and successfully group semantically similar images (some examples of clustering results by TeDeSC are shown in Fig. 1 (top)).
Lastly, we apply TeDeSC to a variety of applications, including discovering biases in real-world datasets and text-to-image generative models, as well as analyzing the popularity of social media images. For example, as shown in Fig. 1(bottom), TeDeSC can uncover less commonly studied biases (e.g., Hair color) in DALL·E3-generated (Betker et al., 2023) images, beyond the well-known biases (e.g., Gender) that may have already been corrected for. It achieves this by discovering various human-interpretable semantic clusters (e.g., “Dark Hair”, “Grey Hair”) across different discovered criteria and identifying overpopulated clusters. Similarly, it can identify the semantic visual elements that contribute to the popularity of social images, offering valuable insights to related practitioners. These results suggest that TeDeSC is an automatic, versatile, and highly practical tool, opening up numerous new application opportunities for future research and providing the potential to generate insights from unstructured visual data at scale across various domains.
2 Semantic Multiple Clustering
In this section, we first formally define the task of Semantic Multiple Clustering (SMC), and then introduce the existing and newly proposed benchmarks for SMC.
Task Definition: Given a collection of $N$ unlabeled images $\mathcal{D}=\{\mathbf{x}_{n}\}_{n=1}^{N}$, the goal of SMC is to categorize $\mathcal{D}$ into semantic clusters based on multiple criteria =$\mathcal{R}=\{R_{l}\}_{l=1}^{L}$, where $L$ is the total number of criteria. In detail, a criterion $R_{l}$ refers to the grouping theme according to which a set of images can be organized (or partitioned), such as Activity, Location, Mood, or Time of Day. All semantic clusters $\mathcal{C}_{l}$ under a given criterion $R_{l}$ should align with the theme of that criterion. For instance, the semantic clusters under the criterion Activity should reflect activities depicted in the images (e.g., “Kayaking”, “Cooking”). The same dataset $\mathcal{D}$, when grouped by a different criterion such as Location, could be clustered into categories like “Outdoor”, “Indoor”, and so on.
[…]
@@ 2 Semantic Multiple Clustering
Figure 2: Overview of the SMC benchmarks. We show all clustering criteria and the corresponding ground-truth labels for the example images. We introduce two new challenging benchmarks: COCO-4c and Food-4c.
Benchmarks: Evaluating SMC methods requires benchmarks that can be partitioned under multiple valid criteria. Currently, only a few benchmarks (Yu et al., 2024) support the evaluation of SMC methods: Fruit-2c (Muresan & Oltean, 2018), Card-2c (Kaggle, 2022), Action-3c (Kwon et al., 2024), and Clevr-4c (Vaze et al., 2024). As shown in Fig. 2, these benchmarks are limited by their object-centric nature with simple backgrounds (e.g., Fruit-2c), an insufficient number of criteria (e.g., up to three in Action-3c), and a lack of photorealism due to synthetic generation (e.g., Clevr-4c).
Given that the data encountered in real-world applications is more complex, we propose two new benchmarks for SMC: Food-4c and COCO-4c. Food-4c is sourced from Food-101 (Bossard et al., 2014), which includes 101 Food type (original annotations), along with new annotations for 15 Cuisine types, 5 Courses types, and 4 Diet preferences, totaling four clustering criteria. Additionally, we constructed COCO-4c using images from COCO-val (Lin et al., 2014), where we annotated four criteria with varying number of clusters: 64 Activity, 19 Location, 20 Mood, and 6 Time of day. Examples of these newly constructed benchmarks are shown in Fig. 2. Further details, including the full list of cluster names and the annotation pipeline, are provided in App. B.
3 Methodology: Text Driven Semantic Multiple Clustering
Partitioning an unstructured image collection into semantic clusters based on different (unknown) criteria is challenging, as it requires reasoning over the visual contents of all the images concurrently. Specifically, the system needs to first find commonalities among the images for discovering the partitioning criterion (or theme) and then group the images into semantic clusters according to the discovered criterion. To tackle the challenging SMC task, we propose a two-stage framework, named Text Driven Semantic Multiple Clustering (TeDeSC), that significantly deviates from the commonly used technique of deep feature-based clustering, and instead uses text descriptions as a proxy to reason over the images and uncover hidden patterns in $\mathcal{D}$. We elaborate the key design differences between TeDeSC and related work in App. A.
System overview: As illustrated in Fig. 3, the proposed TeDeSC consists of two modules: Criteria Proposer and Semantic Grouper. The Criteria Proposer (or Proposer in short) processes the entire image set $\mathcal{D}$ to discover diverse common themes among the images and proposes grouping criteria $\mathcal{R}$ in natural language (e.g., Location). Once the criteria are proposed, the Semantic Grouper (or Grouper in short) discovers distinct semantic clusters that adhere to each criterion $R_{l}$ at varying levels of semantic granularity and assigns images to their respective clusters (e.g., “Climbing gym”). In this work, we explore various design choices for both the Proposer and Grouper, detailed in the following subsections. Full implementation details, including exact prompts, are provided in App. C.
@@ 3 Methodology: Text Driven Semantic Multiple Clustering
Figure 3: TeDeSC is composed of a Criteria Proposer and a Semantic Grouper. (Left) The Proposer processes the entire image set to discover and propose diverse grouping criteria expressed in natural language, which are accumulated in the Criteria Pool. (Middle) The Grouper takes these proposed criteria from the pool, discovers semantic clusters linked to the criteria at three different levels of granularity, and assigns images to their respective clusters. (Right) Multiple substructures of the image set, indicated by the discovered criteria, are revealed across different granularities by combining the cluster assignment results. We explore various design choices, and set the best-performing methods from our experiments as the main configuration (marked with Refer to caption).
@@ 3.1 Criteria Proposer
As shown in Fig. 3(left), the Proposer takes the target image collection as input and outputs a set of discovered grouping criteria, expressed in natural language. The core of the Proposer design lies in its ability to concurrently analyze and reason over the visual content of a large set of images. With this in mind, we explore three systematic approaches, which we describe below.
Image-based Proposer: We start with a baseline that leverages a state-of-the-art MLLM (Li et al., 2024) designed for multi-image reasoning out-of-the-box to directly infer the criteria given a set of images. To reason over a set of images, which is crucial for discovering the criterion, we stitch a batch of images into a $8\times 8$ image grid and input it to the MLLM as a single image. We then prompt the MLLM to propose grouping criteria for the images in the grid. All the resulting criteria from all such image grids are accumulated in a criteria pool, denoted as $\widetilde{\mathcal{R}}$.
Tag-based Proposer: Next, we explore a tag-based approach that uses an open-vocabulary tagger (Radford et al., 2021) to assign 10 tags to each image in $\mathcal{D}$, using the vocabulary from WordNet (Miller, 1995). These tags serve as semantic descriptors, translating the visual content of each image into textual form. We then gather all the assigned tags from the images, input them into a LLM, and prompt it to discover grouping criteria based on these tags.
Caption-based Proposer (Main): While image tags effectively capture certain visual semantics, we found that they predominantly reflect object-related content. To encompass a broader spectrum of visual information—such as environmental settings or interactions—we instead use a MLLM to generate captions for each image in $\mathcal{D}$. Descriptive captions provide a richer and more holistic semantic context. Staying within the 128k token limit of modern LLMs (Meta, 2024b; OpenAI, 2024), we feed a subset of captions into a LLM, which we prompt to elicit partitioning criteria. The criteria generated for each subset are accumulated in $\widetilde{\mathcal{R}}$. Experiments in Sec. 4.1 show that the Caption-based Proposer is the most effective; so we use it as our main method and consider the others as baselines.
Criteria refinement: Since the Proposers operate on subsets, the criteria accumulated in $\widetilde{\mathcal{R}}$ may include redundant or noisy entries. To address this, we input all initially proposed criteria from $\widetilde{\mathcal{R}}$ into a LLM (Meta, 2024b), prompting it to consolidate similar criteria and discard noisy ones. This step refines and updates the criteria pool into $\mathcal{R}$, which is ready to be used in the subsequent stage.

### v4

@@ Abstract
Organizing unstructured image collections into semantic clusters is a long-standing challenge. Traditional deep clustering techniques address this by producing a single data partition, whereas multiple clustering methods uncover diverse alternative partitions, but only when users predefine the clustering criteria. Yet expecting users to specify such criteria a priori for large, unfamiliar datasets is unrealistic. In this work, we introduce the task of Open-ended Semantic Multiple Clustering (OpenSMC), which aims to automatically discover clustering criteria from large, unstructured image collections, revealing interpretable substructures without human input. Our framework, $\bm{\mathcal{X}}$-Cluster: eXploratory Clustering, treats text as a reasoning proxy: it concurrently scans the entire image collection, proposes candidate criteria in natural language, and groups images into meaningful clusters per criterion. To evaluate progress, we release COCO-4c and Food-4c benchmarks, each annotated with four grouping criteria. Experiments show that $\mathcal{X}$-Cluster effectively reveals meaningful partitions and enables downstream applications such as bias discovery and social media image popularity analysis. We will open-source code and data to encourage reproducibility and further research.
Organizing Unstructured Image Collections using Natural Language
Mingxuan Liu1    Zhun Zhong4    Jun Li6    Gianni Franchi3    Subhankar Roy5    Elisa Ricci1,2 1 University of Trento   2 Fondazione Bruno Kessler  3 ENSTA Paris, Institut Polytechnique de Paris 4 Hefei University of Technology   5 University of Bergamo   6 Technical University of Munich
Figure 1: Top: Open-ended Semantic Multiple Clustering (OpenSMC) deals with automatically organizing an unstructured image collection into semantically meaningful and human interpretable clusters, under multiple shared themes or criteria, without requiring any prior knowledge. Bottom: Our system enables various applications like discovering novel biases in text-to-image (T2I) generative models. See our project webpage for code and data: here.
@@ 1 Introduction
When grouping a large, unstructured collection of unlabeled images, several questions arise: i) Which criterion should guide the grouping? ii) Do multiple, equally valid criteria exist? iii) Given a criterion, how can we uncover the corresponding semantic clusters? iv) What should the clusters’ granularity be? Because the answers depend on both the collection’s visual content and the objectives of a specific application, each question is inherently open-ended. Having the ability to automatically resolve them would unlock hidden insights in the data and enable valuable applications, such as social media recommendation systems (Cheng et al., 2024) and dataset auditing (Bianchi et al., 2023). footnotetext: † Corresponding author.
Previous studies have attempted to answer these questions. Deep Clustering (DC) approaches (Caron et al., 2018; Van Gansbeke et al., 2020) iteratively partition an unlabeled dataset but produce only a single partition. Multiple Clustering (MC) methods (Yu et al., 2024), instead, uncover several alternative clustering solutions, offering complementary perspectives. Yet both DC and MC ultimately reflect the inductive biases of the underlying networks, which can deviate from the partitions a user actually needs. Worse, the discovered criteria and clusters, denoted merely by index labels, lack human-interpretable semantics.
To give users explicit control over clustering, Text Criterion conditioned Multiple Clustering (TCMC) (Yao et al., 2024, 2025) approaches let the users specify the desired criterion in natural language e.g., “group playing cards by Suit.” TCMC methods (Kwon et al., 2024; Luo et al., 2024b; Yao et al., 2024, 2025) first employ large language models (LLMs) to generate candidate cluster semantics conditioned on the user-specified criterion, then use vision–language models to assign images to semantically related clusters. Although TCMC methods yield user-guided and interpretable partitions, they presume the user already knows valid grouping criteria for the entire collection. As datasets grow and scenes become more complex (e.g., millions of social-media images (Hu et al., 2014)) knowing such prior knowledge becomes impractical. This leads to the question: Can clustering criteria and semantic substructures (groups) be discovered directly from visual data, without any human input?
We formalize this challenge as the task of Open-ended Semantic Multiple Clustering (OpenSMC). As depicted in Figure 1(top), the objective in OpenSMC is to discover, in natural language, both the clustering criteria and their associated semantic clusters without relying on any human priors. The task is challenging because i) it requires concurrent reasoning over the entire image collection to surface valid criteria, posing a challenge to current vision systems, and ii) it is open-ended as neither user-defined knowledge about the clustering criteria nor clustering granularity is available. Tab. 1 contrasts OpenSMC with existing clustering paradigms. In addition, due to the lack of practical benchmarks with diverse criteria for evaluating the OpenSMC, we introduce two realistic and larger benchmarks–COCO-4c and Food-4c–with annotations for up to four criteria.
Table 1: Comparison of different clustering paradigms. Unlike DC, MC, and TCMC settings, the proposed OpenSMC task does not assume any prior knowledge and offers interpretable results.
DC MC TCMC OpenSMC
Prior Knowledge # Criteria
Text Criteria
Knowledge # Clusters
Output Multiple Clustering
Interpretable
Open-ended
To tackle OpenSMC, we leverage text as a compact proxy for reasoning because current vision models cannot yet encode and reliably reason over large image collections. Consequently, we propose $\bm{\mathcal{X}}$-Cluster: eXploratory Clustering, a two-stage framework that combines multimodal LLMs (MLLMs) (Liu et al., 2024b; Li et al., 2025, 2023a) with text-only LLMs (Meta, 2024a). $\mathcal{X}$-Cluster comprises two modules: a Criteria Proposer, which jointly reasons over all images via their textual representation (embeddings, tags, or captions) to surface valid grouping criteria, and a Semantic Grouper, which uncovers semantic clusters and assigns images to them under each discovered criterion. As shown in Figure 1(top), $\mathcal{X}$-Cluster can help users explore unfamiliar image collections at scale, by automatically grouping the images into semantic clusters (e.g., “Surfing” and “Skateboarding”) under diverse, human-interpretable criteria (e.g., Activity). In this work, we investigate three designs for each module and demonstrate the flexibility and effectiveness of our framework.
Finally, we showcase both the practicality and versatility of $\mathcal{X}$-Cluster across several applications, including uncovering biases in real-world datasets and text-to-image generative models, as well as analyzing the popularity of social media images. For instance, as shown in Figure 1(bottom), the method uncovers well-known biases (e.g., Gender) alongside less explored ones (e.g., Hair color) in DALL·E3 -generated images (Betker et al., 2023). These results confirm that $\mathcal{X}$-Cluster is an automated, practical tool for extracting interpretable insights from large-scale, unstructured visual data, opening new avenues for research and industry.
2 Related Work
Image clustering. Deep clustering learns visual features and produces a single partition of an unlabeled dataset via self-supervision (Caron et al., 2018, 2020; Ren et al., 2024). Multiple clustering extends this idea, seeking multiple non-redundant partitions with data augmentations, diversity losses, or subspace methods (Qi and Davidson, 2009; Hu and Pei, 2018; Yao et al., 2023b; Metaxas et al., 2023; Yu et al., 2024). Despite steady progress, both paradigms share key limitations: i) their results reflect inductive biases of the network and training algorithm, hampering generalization beyond object-centric data and often misaligning with user intent; and ii) clusters are returned as numeric indices rather than human-readable names. In contrast, $\mathcal{X}$-Cluster derives both meaningful criteria and cluster names directly from unlabeled data.
@@ 2 Related Work
While $\mathcal{X}$-Cluster also uses text as its reasoning medium, it departs from TCMC systems in two key aspects: i) it automatically discovers the grouping criteria rather than relying on a user-supplied one; ii) it infers both the number and the names of clusters, without any user input. Together, these features make $\mathcal{X}$-Cluster fully automatic and scalable, and well suited to uncover novel, often overlooked insights in large image collections.
Topic Discovery. OpenSMC is also related to Topic Discovery (Blei et al., 2003) in NLP, which uncovers latent themes in large text corpora. Our work likewise seeks to reveal common thematic concepts, but from visual collections, and does so by using LLMs as the reasoning core.
3 OpenSMC
Task Definition: Given a collection of unlabeled images $\mathcal{D}=\{\mathbf{x}_{n}\}_{n=1}^{N}$, the goal of Open-ended Semantic Multiple Clustering (OpenSMC) is to build a system $\mathcal{H}$ that automatically i) discovers a set of $L$ grouping criteria =$\mathcal{R}=\{R_{l}\}_{l=1}^{L}$ described in natural language, and ii) finds interpretable substructures $\mathcal{O}_{l}$ for each criterion by uncovering semantically meaningful clusters and assigns images to them. Formally, we define an OpenSMC system:
$\mathcal{H}:\mathcal{D}\mapsto\left\{\mathcal{O}_{l}=\Bigl{\{}\mathcal{C}_{k}^{l}=\bigl{(}s_{k}^{l},\mathcal{D}_{k}^{l}\bigr{)}\Bigr{\}}_{k=1}^{K_{l}}\,\middle|\,R_{l}\right\}_{l=1}^{L},$
@@ 3 OpenSMC
An OpenSMC system should find $\mathcal{R}$ and $\mathcal{O}_{l}$ automatically, both expressed in natural language. By contrast, in the TCMC setting humans predefine the criterion $\mathcal{R}$ and the desired cluster count $K$.
@@
Figure 2: OpenSMC benchmarks. We introduce two new challenging benchmarks: COCO-4c and Food-4c. We show all annotated criteria and the corresponding labels for the example images.
Refer to caption(Prompts in App. E.1)(Prompts in App. E.2)
@@
Figure 3: $\bm{\mathcal{X}}$-Cluster overview. (Left) The Criteria Proposer reason over the entire image set via their textual proxies and deposits diverse, natural language grouping criteria in a Criteria Pool. (Middle) The Semantic Grouper draws each criterion from the pool, discovers semantic clusters at three levels of granularity, and assigns images to their proper clusters. (Right) Aggregating these assignments reveals the collection’s semantic substructure across multiple criteria and granularities. We evaluate three design variants for each module and adopt the best-performing configuration (marked Refer to caption) as the default $\mathcal{X}$-Cluster. See App. E for implementation details.
Benchmark: Only a few benchmarks exist that are suitable for evaluating OpenSMC. As shown in Figure 2, current benchmarks: Fruit-2c (Muresan and Oltean, 2018), Card-2c (Kaggle, 2022), Action-3c (Kwon et al., 2024), Clevr-4c (Vaze et al., 2024) are limited because they are object-centric (Fruit-2c), lacking photorealism (Clevr-4c), and too few criteria (e.g., at most three in Action-3c).
To alleviate this issue, we introduce two new benchmarks: Food-4c and COCO-4c. Examples of images are presented in Figure 2. We extended Food-101 dataset (Bossard et al., 2014) to create Food-4c which in total has four grouping criteria. Similarly, we extend COCO-val (Lin et al., 2014) to create COCO-4c that has four grouping criteria. We believe that the newly introduced more realistic benchmarks will play a critical role in evaluating OpenSMC systems. Further benchmark statistics and annotation details are provided in App. C.
4 Method
To recap, the objective of an OpenSMC system is twofold: i) discover meaningful grouping criteria, or themes, in an unstructured image collection, and ii) group the images into semantic clusters for each discovered criterion. This task is particularly challenging because it requires simultaneous reasoning over a large number of images, which the current vision systems are not equipped to handle.
System overview. We propose $\bm{\mathcal{X}}$-Cluster, a two-stage framework that is designed to perform joint reasoning over an entire image collection, as required by the OpenSMC task. At a high level, $\mathcal{X}$-Cluster addresses this task by first translating the visual content of the image collection into a lower-dimensional representation, i.e., natural language descriptions. We then exploit the large context window of the LLMs/MLLMs to discover the grouping criteria, followed by their semantic clusters, using the derived textual descriptions.
Concretely, as shown in Figure 3, $\mathcal{X}$-Cluster comprise: (1) a Criteria Proposer that is in charge of discovering the grouping criteria (e.g., Activity, Mood); and (2) a Semantic Grouper that finds semantic clusters by conditioning the predictions on each discovered criterion (e.g., for a given image “Climbing gym” is predicted when conditioned on the criterion Activity, whereas “Adventurous” is predicted when conditioned on Mood). Next, we elaborate on each component of $\mathcal{X}$-Cluster.
@@ 4.1 Criteria Proposer
We explore three Proposer designs that convert visual information into text, enabling the LLM to infer grouping criteria. Each design differs in how we convert the visual content into text and in how we use the text to derive the grouping criteria.
Image-based Proposer: We prompt a MLLM (Li et al., 2025) to directly infer grouping criteria $\mathcal{R}$ from images, bypassing the translation of visual content into text. In detail, we present a mini-batch of images, tiled into a composite of $8{\times}8$ image grid, to the MLLM. This allows the MLLM to reason concurrently over a group of images. The image-based approach balances broader coverage of criteria with the preservation of fine-grained details, a trade-off introduced by the tiling operation.
Tag-based Proposer. This variant adopts a tag-driven strategy. We use an open-vocabulary tagger (Radford et al., 2021) to assign up to ten WordNet (Miller, 1995) tags to each image, yielding concise textual descriptors. A LLM (Meta, 2024a) then summarizes all the aggregated tags in $\mathcal{D}$ to propose the grouping criteria $\mathcal{R}$.
Caption-based Proposer (Main). While image tags capture certain visual semantics, they primarily focus on object-related content. To incorporate richer visual cues, such as scene context and interactions, we use a MLLM (Liu et al., 2024b) to caption every image in $\mathcal{D}$. We then split the captions into the largest chunks that fit the LLM’s context window (e.g.128k or 1M tokens) (Meta, 2024a; OpenAI, 2024) and prompt the LLM on each chunk to propose grouping criteria. The deduplicated criteria from all chunks form the pool $\mathcal{R}$.